The orbital speed of electron orbiting around the nucleus in a circular orbit of radius 50Pm is 2.2x 106 m/s, the magnetic dipole moment of an electron is
1
Expert's answer
2020-06-18T10:57:57-0400

Given

velocity of an electron is "v=2.2*10^6 m\/s"

Radius of circular path is "r=50*10^{-12} m"

The expression for the magnetic dipole of an electron is

"p=\frac{evr}{2}" "=\frac{(1.60*10^{-19})(2.2*10^6)(50*10^{-12})}{2}=8.80*10^{-24} A.m^2"